OpenStudy (anonymous):

integrate (x+1)/(x^2+1)

OpenStudy (anonymous):

\[\int\limits (x+1)(x^2+1) = \int\limits x ^{3} + x ^{2} + x +1= x^4/4 + x^3/3 + x^2/2 + x\]

OpenStudy (anonymous):

it should be \[(x+1)\div(x^2+1)\]

Directrix (directrix):

.5 ln (x^2+1) + inverse tan (x) + c http://www.wolframalpha.com/input/?i=Integral+of+%5B%28x%2B1%29%2F%28x%5E2%2B1%29%5D+dx